OpenStudy (anonymous):

A city park has an architectural pillar in the shape of a triangular pyramid. The shape of the base of the pillar is an equilateral triangle. The dimensions of the shape of the pillar are as shown below The park plans to repaint the lateral area of the pillar. The area of the pillar that will be repainted is _______________ ft2.

hero (hero):

Those note taking techniques must not be working at all

OpenStudy (anonymous):

@Calcmathlete i absolutely couldnt figure this one out ;/

OpenStudy (anonymous):

I cant use my notes.. its a practice exam, with this practice exxam i can use to complete my final

hero (hero):

What's the point of taking notes if you can't use them? At the very least, you should have memorized the formulas

OpenStudy (anonymous):

I just found out I couldnt use them

OpenStudy (anonymous):

right before starting

OpenStudy (anonymous):

FIrst of all, the lateral area is the surface area of the pyramid EXCLUDING the base. The formula would normally be: SA = (s^2√3)/4 + 3/2sl Lateral Area: LA = 3/2sl where s is a side of the base and l is the slant height.

OpenStudy (anonymous):

so SA =(4^2√3)/4+3/2(4)(11) ?

OpenStudy (anonymous):

72.93?

OpenStudy (anonymous):

Ummm...no. You're finding the lateral area. Use the second formula: LA = 3/2sl where s is a side of the base and l is the slant height.

OpenStudy (anonymous):

la=3/2(11)(4)

OpenStudy (anonymous):

66 inches

OpenStudy (anonymous):

66ft^2

OpenStudy (anonymous):

Yup :) You really just need to remember the formulas and you'd do just fine :)

OpenStudy (anonymous):

thanx ;))

OpenStudy (anonymous):

You're welcome